OpenStudy (anonymous):

the length of a rectangle is 6 less than it's width. the area of the rectangle is 27 cm squared. what is the length and the width of the rectangle to w squared -6 =27 now what?

OpenStudy (anonymous):

l = w - 6 l*w = 27 (w-6)*w = 27 w^2 - 6w - 27 = 0. Solve the quadratic for w. Then keep the positive value. Compute length from there.

OpenStudy (anonymous):

thanks so much sir.